Five years ago Ang was half as old as paula. Now he is 3/5 as old as paula. What is their current ages?
Answer by oberobic(2304)

A = Ang's age
P = Paula's age
.
5 years ago
A-5 = Ang's age
P-5 = Paula's age
.
A-5 = (1/2)*(P-5)
so
A = (1/2)(P-5) +5
.
A = (3/5)*P
.
A=A
so
(1/2)(P-5) +5 = (3/5)*P
.
Multiply both sides by 10 to eliminate the fractions
.
5(P-5) +50 = 6*P
.
5P -25 +50 = 6P
25 = P
P = 25
.
A = 3/5*25
A = 15
.
Answer: Ang is 15, and Paula is 25.
.
Check how the relationship of their ages 5 years ago.
15-5 = 10 = Ang's age then
25-5 = 20 = Paula' age then
Ang's age then was 1/2 Paula's age then.
.
Done.